OpenStudy (aaronandyson):

Find two positive odd number,the sum of whose square is 74. @kidrah69

OpenStudy (kidrah69):

@Nnesha

OpenStudy (kidrah69):

Do you have the equation?

OpenStudy (aaronandyson):

Unable to form it :/

OpenStudy (aaronandyson):

I got 5 & 7

OpenStudy (aaronandyson):

@mathslover help karde yaar

mathslover (mathslover):

Consider two odd numbers: 2n - 1 and 2n + 1 You've : (2n-1)^2 + (2n+1)^2 = 74 Solve for n ;)

OpenStudy (aaronandyson):

yaara thoda simpply method de

mathslover (mathslover):

Kaunsi class mein ho tum?

OpenStudy (aaronandyson):

2 saal k baad yeh chapter karaha hu

mathslover (mathslover):

4n^2 + 1 - 4n + 4n^2 + 1 + 4n = 74 8n^2 + 2 = 74 8n^2 = 72 n^2 = 9 n = +/- 3... since it is given that odd numbers are positive so, n = 0 Therefore, numbers : 2n - 1 = 2(3) - 1 = 5 and 2(3) + 1 =7

OpenStudy (aaronandyson):

muhje yeh equation mila x^2 + 2x - 35 = 0 :/

mathslover (mathslover):

kar lo toh... use middle term splitting method.

OpenStudy (aaronandyson):

(x-5)(x+7) = 0?

mathslover (mathslover):

Yup.

OpenStudy (aaronandyson):

buy its -7 the question says positive kya galt kiya maine??

OpenStudy (aaronandyson):

but*

mathslover (mathslover):

Tumne equation galat nikali hogi

mathslover (mathslover):

how did u come to that equation?

Nnesha (nnesha):

(x-5)(x+7) = 0? are factor form you have to solve for x

OpenStudy (aaronandyson):

mera step 1 yeh tha x^2 + (x + 2)^2 = 74

mathslover (mathslover):

What if x is even?

mathslover (mathslover):

See, you've to get it generalized. odd number has the form of 2n + 1 where n is any integer. So, better use : 2n - 1 and 2n + 1 for ease.

OpenStudy (aaronandyson):

tuh explain kar yar yhe method ni aati

mathslover (mathslover):

You'll have to read my method first. Scroll up and check it out. I've to leave for now for studies. Put in some efforts and you'll get it done.

OpenStudy (aaronandyson):

Chall koi nahi thanks :)

mathslover (mathslover):

And aage se seniors ko respect dena seekho! :) "Chall" .. "Tu" are just too informal to address any one elder to your age. :) Good Luck!

OpenStudy (kidrah69):

ye hui na baat! :D

OpenStudy (aaronandyson):

am a bit tension sorry :(

mathslover (mathslover):

It's alright. Don't get tensed for anything. It will just make your way harder. Find a way to refresh yourself and you'll be done with your worries very soon. Again, good luck!

OpenStudy (aaronandyson):

andplus we had a convo the other day so i though u were of my age

mathslover (mathslover):

:) Remember one thing, it's always good to be formal with the ones who you don't know. Once you start knowing them, you can be a bit informal. And I've no worries with anyone calling me by "Tu" or "Chall" ... It's good! But, I was just suggesting you as keeping in mind about future of yours. This will help you! :)

OpenStudy (aaronandyson):

Thanks for the piece of advice :)

Nnesha (nnesha):

\[\cancel{chall} \]choro na ab sigh lel jk hahahha very true !!! my fvt word is "aap " istead of tu

OpenStudy (aaronandyson):

@Nnesha Can you help me more??

Nnesha (nnesha):

i hve 2 minutes let's see what i can do

OpenStudy (aaronandyson):

Th denominator of a positive number is one less than twice the numinarator .If the sum of their receprocal is 2.9 Find the numbers.

OpenStudy (aaronandyson):

the * numerator* reciprocals * My typos X_X

OpenStudy (aaronandyson):

@Nnesha